# Envisioning New Futures of Positive Social Technology: Beyond Paradigms of Fixing, Protecting, and Preventing

### v1

@@ 1. Introduction
Much of social computing research today primarily focuses on addressing the flaws of current platforms, particularly in social media. However, this often limits exploring new, potentially more beneficial socio-technical systems. Indeed, today’s research interventions are often confined within the bounds of platforms like Facebook and Instagram (González-Bailón et al., 2023; Guess et al., 2023a, b; Nyhan et al., 2023), as opposed to pioneering new models, such as in research prototypes from two decades ago (Miller et al., 2003). Furthermore, legislative approaches focus reactively on addressing problems on specific, existing platforms rather than proactively considering new possibilities  (pre, [n.d.]; Schatz and [d Hi], 2023; Stewart and [r Ut-2], 2023; bil, [n.d.]) for the future of social technologies, furthering the view of social media as a fixed entity.
@@ 1.1. Positive Social Technology (Positech)
Figure 1. Venn Diagram of Positech and Related Concepts
[…]
The image depicts a Venn diagram with three overlapping circles within a larger rectangular frame titled “Positech.” Each circle represents a different field: “Positive Technology,” “Positive Computing,” and “Asset-based Approach.” The overlapping sections of the circles are labeled with letters (A), (B), (C), (D), and (E) to indicate areas of intersection between these fields. Area (A) is the portion of the “Positive Technology” circle that does not overlap with the other fields. Area (B) is the portion of the “Positive Computing” circle that does not overlap with the other fields. Area (C) is the portion of the “Asset-based Approach” circle that does not overlap with the other fields and is outside of the “Positech” rectangle. Area (D) is the portion within Positech that is outside the all other circles, indicating the unique portion of the “Positech” rectangle. Area (E) is the central overlapping area where all three circles intersect, indicating a shared space between “Positive Technology,” “Positive Computing,” and “Asset-based Approach.”
Figure 1 shows the relationships between Positech and its adjacent fields. Here, we lay out the distinct guiding questions of each field for each labeled area in the figure.
@@
An example outcome at the intersection of these fields (labeled (E) in Figure 1) could include research on a proof-of-concept for a new social media platform designed to explore new possibilities for increasing human connectedness. This platform would leverage people’s intrinsic motivations to empathize with others to enhance the effectiveness of peer support.
@@ 2. Organizers
JaeWon Kim is a PhD candidate at the University of Washington Information School. Her research focuses on understanding how social technologies support or undermine relationship building and designing and building social technologies that center on trusting relationship building, especially for the youth.
@@
3. Agenda
Our workshop at CSCW 2024 will bring together 15-30 researchers/ practitioners/designers to discuss the area of Positech: aiming to envision new futures of positive social technology and strategize how to support each other’s efforts. While we acknowledge that the absence of a remote participation option may present equity concerns for some potential participants, we will be holding this workshop in person only due to budgetary and logistical limitations.
We aim to explore the overlapping fields of research that contribute to Positech, project a current landscape and future agenda, and plan how to build community and support our peers. This workshop puts a specific focus on community-building and collaborative agenda-setting; we, therefore, dedicate most of our time to small-group brainstorms and discussions. We also plan to collaboratively develop a course syllabus through the workshop to advance and project the Positech research agenda.
Table 1. Workshop Agenda
Time Activity
9:00-10:00am Welcome and Attendee Introduction
10:00-10:20am Coffee Break
10:20-11:20am Panel: Positech Histories and Influences
11:20-12:00pm Discussion: Positech Boundaries and Overlaps
12:00-1:30pm Lunch Break
1:30-2:50pm Brainstorm, Share, Synthesize: Positech Challenges and Promising Directions
2:50-3:10pm Coffee Break
3:10-3:50pm Brainstorm, Share: How to Support Positech Research
3:50-4:40pm Brainstorm, Share: Positech Community Futures
4:40-5:00pm Closing
6:30pm (Optional) Dinner Social Event
3.1. Themes
Our workshop activities are motivated by particular goals, which we realize via three themes.
3.1.1. Theme 1: Exploring Relevant and Adjacent Research
We explore this theme via the “Positech Histories and Influences” Panel. We will invite panelists across fields that overlap with Positech to explore the following questions:
  • Historical Lessons: What can we learn from prior research on social technologies? How can that inform future research directions? How have previous platforms/research projects influenced current social technologies?
  • Defining Positech: What do we think are important distinctions to draw between Positech and related fields or frameworks? How can we synthesize insights from related fields such as positive computing, positive technologies, strength-based technology, and asset-based technology with lessons from adjacent disciplines like psychology to inform the development of social technologies aimed at human flourishing?
3.1.2. Theme 2: Projecting the Landscape of Positech
We address this theme via two Discussions: “Positech Boundaries and Overlaps” and “Positech Challenges and Promising Directions.”
We aim to map the current landscape of Positech. Key questions for discussion include: What constitutes Positech? What are its challenges and opportunities? To facilitate this exploration, participants will collaborate in groups to design a course syllabus regarding themes around Positech: origins, overlapping fields and frameworks, and different ongoing research directions.
3.1.3. Theme 3: Envisioning the Future of Positech
This theme is addressed through the “How to Support Positech Research” and "Positech Community Futures” Discussions.
For the final activity of our workshop, we will engage in a brainstorming session to outline our post-workshop plans. This will be conceptualized as a research question: How can we maintain an asynchronous, geographically dispersed, cross-generational, and (potentially) culturally diverse research community? Addressing community continuity is a complex challenge that has garnered much attention in social computing  (Bali et al., 2023; Burke et al., 2009; Kraut et al., 2011), and social technologies are perhaps our best tool for sustaining such a community across geographical distances.  [^1] One example idea that may come out of this discussion is a Slack  (Slack, [n.d.]) bot designed to facilitate interaction among participants. This bot would pair members for virtual coffee chats, using conversation starters linked to their research interests to foster sustained dialogue and collaboration.
We plan to implement strategies developed during the session to ensure the momentum generated by our workshop continues to build. Relevant questions include:
  • Ideal scenarios. What does an ideal future look like with Positech fully integrated into social technology research and society in general?
  • Sustaining Positech. How can we keep the momentum alive post-workshop? How should this community function and interact? How can we assist others in the Positech community with research?
  • Unintended consequences. How can we anticipate unintended consequences of new technologies or communities proactively and prevent them at an early stage? How can we ensure that Positech actually prioritizes and supports the well-being of individuals and communities?
3.2. Deliverables
We aim to come away from the workshop with two main takeaways:
  1. (1)
    Positech Syllabus: a course syllabus on Positech. This syllabus will outline the course’s overview and motivation, feature 8-15 weekly themes covering various Positech aspects, include recommended weekly readings, and detail assignments such as projects and reading responses. We will share the result on our website as a resource for the wider community, along with a broader list of exemplary papers. (Developed during first two discussions, including insights from the panel.)
  2. (2)
    Positech Community Plan: a set of steps to develop and maintain a community of researchers engaged in the Positech vision. This will include ideas for maintaining communication between attendees, facilitating discovery for newcomers to the area, and ways to connect members with support within the community. (Developed during last two discussions.)
4. Workshop Logistics
4.1. Website
The link to our website is https://positech-cscw-2024.github.io/. We have prepared the website with an overview of our workshop, a call for participation, and key information about the workshop, including the workshop agenda and organizer information. We also plan to publicize any outcome of our workshop, such as the hypothetical Positech course syllabi and workshop submissions, if given the authors’ consent.
4.2. Pre-workshop Plan
Call for participation. We will invite 25-30 individuals interested in the goal of envisioning new futures of positive social technology to participate in our in-person-format workshop. Submissions are welcomed in several forms:
  • Position papers or drafts (2-6 pages, ACM single-column format, excluding references) discussing themes of the workshop. Alternate forms, such as design fiction, are encouraged.
  • “Encore” submissions of relevant conference or journal papers.
  • Statements of research interest (up to 2 pages, ACM single-column format, excluding references).
Selecting participants. Submissions will be reviewed by the organizers based on: 1) relevance to the workshop’s theme, 2) quality of the submission, and 3) diversity of perspectives. This approach will ensure a rich exchange of ideas among participants from varied backgrounds and specialties.
Workshop preparation. Prior to the workshop, we will survey participants to understand their accessibility needs, constraints, and expectations. This feedback will inform the logistics and schedule, with experienced organizers ensuring an engaging experience for both in-person and remote attendees.
4.3. Accessibility
We will ask all attendees to supply an accessable PDF of their submissions. For online workshop attendees, we will enable (auto-generated) live closed captioning, which participants will be able to turn on as desired. Any special accessibility requirements can be communicated to us, and we will do our best to provide an accommodation in partnership with the Accessibility Chair.
4.4. Post-workshop plan
Our workshop is designed to generate (a) a syllabus, and (b) a plan for community building after the workshop. We plan to make the syllabus widely available and publicize it after the workshop. We also plan to follow the community-building plan to keep our ties and build new ones after the workshop.

### v3

@@ 1. Introduction
Much of social computing research today primarily focuses on addressing the flaws of current platforms, particularly in social media. However, this often limits exploring new, potentially more beneficial socio-technical systems. Indeed, today’s research interventions are often confined within the bounds of platforms like Facebook and Instagram (González-Bailón et al., 2023; Guess et al., 2023a, b; Nyhan et al., 2023), as opposed to pioneering new models, such as in research prototypes from two decades ago (Miller et al., 2003). Furthermore, legislative approaches focus reactively on addressing problems on specific, existing platforms rather than proactively considering new possibilities  (Skinner, [n. d.]; CONGRESS.GOV, 2023a, b; of State Legislatures, NCSL) for the future of social technologies, furthering the view of social media as a fixed entity.
@@ 1.1. Positive Social Technology (Positech)
Here, we lay out the distinct guiding questions of each adjacent field to Positech.
@@
An example outcome at the intersection of these fields could include research on a proof-of-concept for a new social media platform designed to explore new possibilities for increasing human connectedness. This platform would leverage people’s intrinsic motivations to empathize with others to enhance the effectiveness of peer support.
@@ 2. Organizers
JaeWon Kim is a PhD candidate at the University of Washington Information School. Her research focuses on understanding, designing, and building social technologies that center on meaningful social connections, especially for the youth.
@@
3. Workshop Logistics and Agenda
Our workshop at CSCW 2024 will bring together 15-30 researchers /practitioners/designers to discuss the area of Positech: aiming to envision new futures of positive social technology and strategize how to support each other’s efforts. We aim to explore the overlapping fields of research that contribute to Positech, project a current landscape and future agenda, and plan how to build community and support our peers. More information on the agenda can be found on our workshop website: https://positech-cscw-2024.github.io/. We plan to publicize any outcome of our workshop.